Astronauts on the first trip to Mars take along a pendulum that


has a period on earth of 1.50 s. The period on Mars turns out to be


2.45 s. What is the free-fall acceleration on Mars?

1
Expert's answer
2021-12-23T10:48:39-0500

Solution:



"T=2\pi \sqrt{\frac{l}{g}}"

g at earth is:


"g=9.81\frac{m}{s^2}"


"\frac{T_{mars}}{T_{earth}}=\sqrt{\frac{g_{earth}}{g_{mars}}}"


"\implies \frac{2.45}{1.50}=\sqrt{\frac{9.81}{g_{mars}}}\implies g_{mars}=3.677\frac{m}{s^2}"


Answer:


"g_{mars}=3.677\frac{m}{s^2}"